# Bayesianize Fuzziness in the Statistical Analysis of Fuzzy Data

### v1

@@ 6.4 Case study
The approximated Gibbs sampler described in Section 4 was used on the final dataset, with five independent chains of 4000 samples each (burn-in period: 2000 samples). As for the previous applications, weakly informative priors were used. Since the response variable consists of rating data, the $(1,5)$-Beta distribution with the mean-precision parametrization was used to model the non-fuzzy component, namely $f_{Y}(y;\bm{\theta}_{y})=\mathcal{B}e(y;\mu\phi,\phi-\mu\phi)_{[1,5]}$. The variables age, rel_length, sex_desire, gender_partner, and respo_partner were used in the linear predictor $\eta_{i}$ for $\mu_{i}=\mathbb{E}\left[Y_{i}\right]=g^{-1}(\eta_{i})$ (as usual, the logistic function was used as invere link function $g^{-1}$). All the chains reached the covergence according to the Gelman and Rubin’s $\widehat{R}$ index and were subsequently used in the data analysis. Table 3 reports some of the posterior statistics, the 95% Highest Posterior Density Interval (HPDI) bounds, and the Effective Sample Size (ESS) diagnostics. Figure 7 shows the marginal posterior densities for the model parameters.
@@ 7 Conclusions
This paper introduces a conditional probabilistic framework for the statistical analysis of bounded fuzzy data, providing a consistent and comprehensive approach to handling fuzzy variables. While numerous fuzzy statistical tools exist, they often lack a unified framework integrating general, widely applicable parametric probability models for fuzzy random variables. This gap has hindered their practical application, particularly in delivering broadly applicable inferential results. Building on the premise that fuzzy data arise from a process akin to statistical coarsening, we propose a regression-based approach that links observed fuzzy statistics – such as modes and spreads – to their underlying statistical model counterparts. This model also accommodates covariates within a more general approximated Gibbs sampling-based solution, enhancing its flexibility and applicability. Through simulations and case studies, we demonstrate the effectiveness of the proposed method, showcasing its ability to incorporate fuzziness within a structured inferential setting. Furthermore, leveraging a Bayesian approach enhances both the interpretability and inferential capacity of fuzzy data analysis, making it a powerful tool for statistical inference in uncertain environments.
@@
An additional contribution of this paper pertains to the ongoing debate on the simulation of fuzzy numbers. While traditional methods have predominantly relied on deterministic or arithmetic approaches for generating fuzzy numbers (see [23, 25, 24] for a detailed discussion), our approach adopts a probabilistic framework grounded in Bayesian methods. This integration embeds the simulation process within a broader inferential paradigm, offering a unified statistical representation that seamlessly handles both data simulation and analysis. In particular, our method aligns naturally with the parametric regression framework, enhancing its applicability in modeling and inference.
@@
Despite its strengths, our approach is not without limitations. Notably, it reduces fuzziness to a collection of summary statistics—albeit meaningful ones—potentially sacrificing some of its full expressive power. Additionally, the proposed method does not incorporate an integrated possibilistic framework for statistical inference, which could better capture fuzzy data’s inherent richness and complexity. However, it is essential to recognize that the scientific community currently lacks a fully possibilistic approach that matches the expressiveness and inferential capabilities of probabilistically grounded statistical frameworks. This trade-off, though common in statistical modeling, parallels challenges in other fields, such as network data analysis, where simplifications are often made to enable inference. This raises a fundamental question: To what extent does replacing fuzzy numbers with summary statistics represent real progress, and could this shift be more meaningfully interpreted within the broader statistical paradigm? After six decades of research in fuzzy set theory, it is worth reconsidering the role of fuzzy statistics in modeling imprecise data and whether its goals can be effectively incorporated into existing statistical approaches (e.g., Bayesian Networks). The challenge of addressing and ultimately resolving this question falls to future generations of scholars.

### v2

@@ 6.4 Case study
The approximated Gibbs sampler described in Section 4 was used on the final dataset, with five independent chains of 4000 samples each (burn-in period: 2000 samples). As for the previous applications, weakly informative priors were used. Since the response variable consists of rating data, the $(1,5)$-Beta distribution with the mean-precision parametrization was used to model the non-fuzzy component, namely $f_{Y}(y;\bm{\theta}_{y})=\mathcal{B}e(y;\mu\phi,\phi-\mu\phi)_{[1,5]}$. The variables age, rel_length, sex_desire, gender_partner, and respo_partner were used in the linear predictor $\eta_{i}$ for $\mu_{i}=\mathbb{E}\left[Y_{i}\right]=g^{-1}(\eta_{i})$ (as usual, the logistic function was used as invere link function $g^{-1}$). All the chains reached the convergence according to the Gelman and Rubin’s $\widehat{R}$ index and were subsequently used in the data analysis. Table 3 reports some of the posterior statistics, the 95% Highest Posterior Density Interval (HPDI) bounds, and the Effective Sample Size (ESS) diagnostics. Figure 7 shows the marginal posterior densities for the model parameters.
@@
The Supplementary Materials (Sections 2.1-2.2) contain extended numerical and graphical results for this case study, including a comprehensive evaluation of the results obtained using the proposed conditional model compared to those from a selection of state-of-the-art regression methods for fuzzy data.
@@ 7 Conclusions
This paper introduces a conditional probabilistic framework for the statistical analysis of bounded fuzzy data, providing a consistent and comprehensive approach to handling fuzzy variables. While numerous fuzzy statistical tools exist, they often lack a unified framework integrating general, widely applicable parametric probability models for fuzzy random variables. This gap has hindered their practical application, particularly in delivering broadly applicable inferential results. Building on the premise that fuzzy data arise from a process akin to statistical coarsening, we propose a regression-based approach that links observed fuzzy statistics – such as modes and spreads – to their underlying statistical model counterparts. This model also accommodates covariates within a more general approximated Gibbs sampling-based solution, enhancing its flexibility and applicability. Through simulations and case studies, we illustrate the effectiveness of the proposed method, showcasing its ability to incorporate fuzziness within a structured inferential setting. Furthermore, leveraging a Bayesian approach enhances both the interpretability and inferential capacity of fuzzy data analysis.
Significantly, the conditional approach proposed in this work is grounded in an epistemic interpretation of fuzziness, in which observed fuzzy values are treated as imprecise representations of latent, unobserved random variables. This assumption does not rely on any specific cognitive mechanism on the part of the observer; instead, it provides a statistically principled way to model non-stochastic imprecision. Consequently, the proposed framework is applicable across multiple domains, such as sensory evaluations, questionnaire-based ratings, and decision support systems, where individuals often provide fuzzy assessments due to limited knowledge or intrinsic measurement imprecision. In all these cases, the blurring mechanism serves to approximate how non-fuzzy random quantities are transformed into fuzzy responses through the measurement process.
@@
An additional contribution of this paper pertains to the ongoing debate on the simulation of fuzzy numbers. While traditional methods have predominantly relied on deterministic or arithmetic approaches for generating fuzzy numbers (see [25, 27, 26] for a detailed discussion), our approach adopts a probabilistic framework grounded in Bayesian methods. This integration embeds the simulation process within a broader inferential paradigm, offering a unified statistical representation that seamlessly handles data simulation and analysis. In particular, our method aligns naturally with the parametric regression framework, enhancing its applicability in modeling and inference. Other epistemic sampling processes considered previously in the literature [28, 25, 8, 37] are usually more straightforward in their nature, albeit without the significant “Bayes flavor” existing in the proposed approach (9)–(11).
By and large, this work emphasizes that, as in other areas of mathematical and statistical modeling, a solid foundation for the proposed constructions remains essential. In particular, we think that any data analysis with fuzzy numbers need to be properly formulated within a probabilistic framework, which ensures that the methodology developed for analyzing imprecise data based on random fuzzy numbers is statistically well-justified. While many of the solutions available in the literature on fuzzy data analysis are largely heuristic, the approaches proposed in this work are firmly grounded in the methodology of standard statistical inference. Special attention is given to the beta-type fuzzy numbers considered here, which, in our opinion, deserve particular emphasis due to their ease of use, interpretation, generation, understanding, and visualization, as well as their convenience for Bayesian analysis considering that most of the existing literature restricts itself to trapezoidal fuzzy numbers.
@@
Despite its strengths, our approach is not without limitations. Notably, it reduces fuzziness to a collection of summary statistics – albeit meaningful ones – potentially sacrificing some of its full expressive power. Additionally, the proposed method does not incorporate an integrated possibilistic framework for statistical inference, which could better capture fuzzy data’s inherent richness and complexity. However, it is essential to recognize that the scientific community currently lacks a fully possibilistic approach that matches the expressiveness and inferential capabilities of probabilistically grounded statistical frameworks. This trade-off, though common in statistical modeling, parallels challenges in other fields, such as network data analysis, where simplifications are often made to enable inference. This raises a fundamental question: To what extent does replacing fuzzy numbers with summary statistics represent real progress, and could this shift be more meaningfully interpreted within the broader statistical paradigm? After six decades of research in fuzzy set theory, it is worth reconsidering the role of fuzzy statistics in modeling imprecise data and whether its goals can be effectively incorporated into existing statistical approaches (e.g., Bayesian Networks). The challenge of addressing and ultimately resolving this question falls to future generations of scholars.